3)1200 soldiers in a fort had enough food for 28 days. After 4 days, some soldiers were sent to another fort and thus, the food lasted for 32 more days. How many soldiers left the fort?
Answer by Narsinghchauhan8960103366(17)

Let x soldiers were transferred to another fort
Remaining soldiers = (1200-x)
Remaining days = (28-4)=24 days
According to question
1200/(1200-x)=32/24
1200*24=(1200-x)*32
28800=38400-32x
32x=38400-28800=9600
x=9600/32=300
So transferred soldiers 300 ans